# Localized statistics decoding: A parallel decoding algorithm for quantum low-density parity-check codes

### v1

@@ Appendix B Parallel implementation
In this section, we propose a parallel version of LSD and analyze its time complexity. The central observations are that clusters can be grown and solved (using the on-the-fly elimination technique outlined in Section 4.2) in parallel and hence potential synchronization issues arise only due to cluster merges.
To enable efficient parallel merging of clusters, we represent the clusters by a parallel version of the union-find data structure as proposed in Ref. [58], where the authors prove that the union-find operations can be performed in parallel polylogarithmic time on each parallel resource with such a data structure. We detail the key steps of the parallel implementation of the LSD algorithm, P-LSD, in Algorithm 2.
@@
    (Parallel) initialization: we create a cluster $C_{i}$ for each flipped detector node $s_{i}$ in the syndrome. We then compute the set of candidate fault nodes at the boundary of each cluster: $\Lambda(C_{i})=\Gamma(\{s_{i}\})$ – see Eq. 4. Clusters are managed as a union-find forest in the parallel data structure $\mathbb{U}$, see Ref. [65].
@@
    (Parallel + synchronization) growth step: add one new fault node to each cluster and merge colliding clusters (Definition 4.4) using the parallel union-find data structure $\mathbb{U}$ and its efficient union-find operations. The parallel_union routine [58] is used to efficiently determine which set in $\mathbb{U}$ a node belongs to and to efficiently merge any two sets in $\mathbb{U}$ that correspond to colliding clusters.
@@
    (Parallel) validity check: Apply on-the-fly elimination in parallel for each cluster, see Section 4.2.
@@ 4 $\mathbb{U}$: parallel union-find data structure
6       $\mathbb{U}$leftarrow initialize_cluster($s_{i}$)
@@
    Pre-growth step: iterate over the candidate fault nodes for each cluster, and identify the most suitable fault node given the soft information vector $\mathbf{\lambda}$. This step can be performed in logarithmic complexity in the size of the boundary using, e.g., a Fibonacci heap. Maintaining a list of candidate fault nodes (Eq. 4) for each cluster throughout the algorithm incurs only a constant depth overhead.
@@
In conclusion, the overall parallel depth for an iteration of the while loop in Algorithm 2 of Algorithm 2 is in $O(\operatorname{polylog}(n)+\kappa^{3})$, where $\kappa$ is the maximum cluster size. Crucially, the maximum cluster size $\kappa$ is expected to be small for QLDPC codes and low enough error rates, see Appendix A. In fact, the average cluster size $\kappa_{\alpha}$ can be bounded for regular graphs, such as the Bethe lattice by $\kappa_{\alpha}\leq\frac{1+p}{1-(\theta-1)p}$, where $\theta$ is the vertex degree of the graph and $p$ is the physical noise strength.
The $O(\operatorname{polylog}(n)+\kappa^{3})$ depth bound is a loose upper bound and in practice the runtime of the algorithm depends on multiple additional factors such as the concrete implementation, overhead of parallelization, and the number of parallel resources available. The overall work, i.e., the total number of steps the algorithm performs across all processes depends on the number of growth steps. In the worst case, the number of growth steps is proportional to the number of edges in the decoding graph, and the parallel union-find data structure will have overall work almost linear in $n$, more precisely $O\left(n\alpha(n)\right)$, where $\alpha(\cdot)$ is the inverse Ackerman function, which is $\alpha(n)\leq 3$ for any practical situation [59, 58, 66]. However, if almost $n$ growth steps were necessary for P-LSD to terminate, the computed error estimate would probably have weight above the percolation threshold of the graph, in which case decoding introduces a non-trivial logical error with high probability. In other words, we expect almost $n$ growth steps only in instances of the decoding problem that have an intrinsically non-local structure, meaning instances that can not be divided into smaller decoding problems. In such a case, there is no benefit in parallelization and hence the parallel runtime complexity is only marginally meaningful.
@@ C.1 Surface code decoding
The BP+UF algorithm [37] is a combination of BP and Union-Find, where the BP soft information is used to guide the cluster growth for the Union-Find algorithm. Note that in the case of a matchable code, the validity check for clusters, which is equivalent to solving systems of linear equations in general, can be replaced by a simple parity computation of the support vector of enclosed, flipped detector nodes and hence is more efficient. Moreover, computing a solution for the Union-Find clusters can be done using a simple algorithm based on spanning-tree construction and thus is also efficient for matchable codes. Note that BP+UF and BP+LSD share an identical cluster growth strategy and therefore should converge to the same set of valid clusters, with the difference that BP+UF applies the peeling decoder [59] for finding a valid correction for each cluster, while the LSD algorithm performs a (partial) inversion of the cluster check matrices $H_{[C]}$ to determine the correction.
@@ C.2 Hypergraph product codes
We investigate the decoding performance of a family of hypergraph product codes [41] with rate $k/n\geq 1/25$. The codes are single-shot decodable [69, 37], that is, a constant-sized decoding window is sufficient for decoding codes of arbitrary size. This family of codes has been previously proposed and investigated in Ref. [70] under a code capacity noise model, and later in Ref. [43] under a circuit-level noise model. In both cases, decoding was achieved by a combination of BP and small set flip (SSF) [10]. Additionally, in Ref. [13] the authors investigated this code family under a phenomenological noise model derived from a circuit-level noise model that is inspired by a potential implementation in neutral atom arrays using BP for the bulk of the decoding problem and BP+OSD in the last decoding round to ensure a correction that projects state back into the code space. This family of codes is obtained from random $(3,4)-$regular Tanner graphs with girth at least 6. Since the construction is random, following Ref. [43], we generate 100 instances of each check matrix and select the best one after performing code capacity simulations. Interestingly, we find that in some cases the performance difference between the best and the worst performing codes is close to a factor of 10. Here, we investigate the sub-threshold performance for a circuit-level noise model for 4 instances of the family using our proposed BP+LSD decoder.
@@ C.3 Bivariate bicycle codes
In Figure 6, we show the logical $Z$ error rate per syndrome cycle $p_{L_{Z}}$ for the BB code family decoded with BP+LSD-0. Our results show that BP+LSD-0 achieves comparable decoding performance to the BP+OSD-CS-7 decoding results presented in Ref. [12] (where BP+OSD-CS-7 refers to the “combination sweep” strategy for OSD higher-order processing with order $w=7$, see Ref. [21]). This implies that, at least for the case of the BB code family, higher-order reprocessing has minimal impact when decoding in the sub-threshold regime. See Appendix D for further discussion of higher-order reprocessing.
Another interesting observation from our numerical decoding simulations of BB codes is that it is sufficient to run BP+LSD with a low number of BP iterations: the BP+LSD simulations shown in Figure 6 were run with maximum $30$ BP iterations, compared to the maximum number of iterations of $10,000$ for the BP+OSD-CS-7 simulations in Ref. [12]. For BP-OSD, it is preferable to run BP until convergence to avoid the cost of computing the inverse of the detector check matrix. In OSD, matrix inversion is performed globally with worst-case cubic runtime complexity in the number of fault nodes in the Tanner graph $n$. Therefore, the OSD runtime does not scale with the error rate and is independent of the number of flipped detectors. In contrast, this is not the case for our proposed BP+LSD decoder. When run in parallel, the LSD algorithm has an expected runtime in $O(\kappa^{3})$, where $\kappa$ is the maximum cluster size. Further to this, the number of clusters upon initialization is identical to the number of invalid detectors. We therefore expect the decoding time to be proportional to the physical error rate. As such, in the sub-threshold regime, LSD post-processing is much less costly than BP+OSD. As a result, it is not detrimental to the decoder’s runtime if a large portion of the computational load is shouldered by the LSD post-processing.
@@
[…]
@@ Appendix D Higher-order LSD
The central idea of higher-order LSD is to add an additional reprocessing step to LSD00-0- 0 to achieve better decoding accuracy. Once LSD00-0- 0 has terminated, meaning that all clusters are valid and local solutions have been computed, we conduct, up to $\mu$ additional growth step on each valid cluster. On the grown clusters, we then apply standard OSD reprocessing: we do this separately on each cluster matrix and therefore we preserve LSD locality. We note that:

### v2

@@ Appendix B Parallel implementation
In this section, we propose a parallel version of LSD and analyze its time complexity. The central observations are that clusters can be grown and solved (using the on-the-fly elimination technique outlined in Methods 4.2) in parallel and hence potential synchronization issues arise only due to cluster merges.
To enable efficient parallel merging of clusters, we represent the clusters by a parallel version of the union-find data structure as proposed in Ref. [60], where the authors prove that the union-find operations can be performed in parallel polylogarithmic time on each parallel resource with such a data structure. We detail the key steps of the parallel implementation of the LSD algorithm, P-LSD, in Algorithm˜2.
@@
    (Parallel) initialization: we create a cluster $C_{i}$ for each flipped detector node $s_{i}$ in the syndrome. We then compute the set of candidate fault nodes at the boundary of each cluster: $\Lambda(C_{i})=\Gamma(\{s_{i}\})$ – see Def 4.3. Clusters are managed as a union-find forest in the parallel data structure $\mathbb{U}$, see Ref. [67].
@@
    (Parallel + synchronization) growth step: add one new fault node to each cluster and merge colliding clusters [Def. 4.4 (Cluster collisions)] using the parallel union-find data structure $\mathbb{U}$ and its efficient union-find operations. The parallel_union routine [60] is used to efficiently determine which set in $\mathbb{U}$ a node belongs to and to efficiently merge any two sets in $\mathbb{U}$ that correspond to colliding clusters.
@@
    (Parallel) validity check: Apply on-the-fly elimination in parallel for each cluster, see Methods 4.2..
@@ 4 $\mathbb{U}$: parallel union-find data structure
6 $\mathbb{U}$ $\leftarrow$ initialize_cluster($s_{i}$)
@@
    Pre-growth step: iterate over the candidate fault nodes for each cluster, and identify the most suitable fault node given the soft information vector $\mathbf{\lambda}$. This step can be performed in logarithmic complexity in the size of the boundary using, e.g., a Fibonacci heap. Maintaining a list of candidate fault nodes [Def. 4.3 (Cluster-boundary] for each cluster throughout the algorithm incurs only a constant depth overhead.
@@
In conclusion, the overall parallel depth for an iteration of the while loop in Algorithm˜2 of Algorithm˜2 is in $O(\operatorname{polylog}(n)+\kappa^{3})$, where $\kappa$ is the maximum cluster size. Crucially, the maximum cluster size $\kappa$ is expected to be small for QLDPC codes and low enough error rates, see Appendix˜A. In fact, the average cluster size $\kappa_{\alpha}$ can be bounded for regular graphs, such as the Bethe lattice by $\kappa_{\alpha}\leq\frac{1+p}{1-(\theta-1)p}$, where $\theta$ is the vertex degree of the graph and $p$ is the physical noise strength.
The $O(\operatorname{polylog}(n)+\kappa^{3})$ depth bound is a loose upper bound and in practice the runtime of the algorithm depends on multiple additional factors such as the concrete implementation, overhead of parallelization, and the number of parallel resources available. The overall work, i.e., the total number of steps the algorithm performs across all processes, depends on the number of growth steps. In the worst case, the number of growth steps is proportional to the number of edges in the decoding graph, and the parallel union-find data structure will have overall work almost linear in $n$, more precisely $O\left(n\alpha(n)\right)$, where $\alpha(\cdot)$ is the inverse Ackerman function, which is $\alpha(n)\leq 3$ for any practical situation [61, 60, 68]. However, if almost $n$ growth steps were necessary for P-LSD to terminate, the computed error estimate would probably have weight above the percolation threshold of the graph, in which case decoding introduces a non-trivial logical error with high probability. In other words, we expect almost $n$ growth steps only in instances of the decoding problem that have an intrinsically non-local structure, meaning instances that can not be divided into smaller decoding problems. In such a case, there is no benefit in parallelization and hence the parallel runtime complexity is only marginally meaningful.
@@ C.1 Surface code decoding
The BP+UF algorithm [38] is a combination of BP and Union-Find, where the BP soft information is used to guide the cluster growth for the Union-Find algorithm. Note that in the case of a matchable code, the validity check for clusters, which is equivalent to solving systems of linear equations in general, can be replaced by a simple parity computation of the support vector of enclosed, flipped detector nodes and hence is more efficient. Moreover, computing a solution for the Union-Find clusters can be done using a simple algorithm based on spanning-tree construction and thus is also efficient for matchable codes. Note that BP+UF and BP+LSD share an identical cluster growth strategy and therefore should converge to the same set of valid clusters, with the difference that BP+UF applies the peeling decoder [61] for finding a valid correction for each cluster, while the LSD algorithm performs a (partial) inversion of the cluster check matrices $H_{[C]}$ to determine the correction.x
@@ C.2 Hypergraph product codes
We investigate the decoding performance of a family of hypergraph product (HGP) codes [43] with rate $k/n\geq 1/25$. The codes are single-shot decodable [71, 38], that is, a constant-sized decoding window is sufficient for decoding codes of arbitrary size. This family of codes has been previously proposed and investigated in Ref. [72] under a code capacity noise model, and later in Ref. [45] under a circuit-level noise model. In both cases, decoding was achieved by a combination of BP and small set flip (SSF) [10]. Additionally, in Ref. [13] the authors investigated this code family under a phenomenological noise model derived from a circuit-level noise model that is inspired by a potential implementation in neutral atom arrays using BP for the bulk of the decoding problem and BP+OSD in the last decoding round to ensure a correction that projects state back into the code space. This family of codes is obtained from random $(3,4)$-regular Tanner graphs with girth at least 6. Since the construction is random, following Ref. [45], we generate 100 instances of each check matrix and select the best one after performing code capacity simulations. Interestingly, we find that in some cases the performance difference between the best and the worst performing codes is close to a factor of 10. Here, we investigate the sub-threshold performance for a circuit-level noise model for 4 instances of the family using our proposed BP+LSD decoder.
@@ C.3 Bivariate bicycle codes
In Fig. 6, we show the logical $Z$ error rate per syndrome cycle $p_{L_{Z}}$ for the BB code family decoded with BP+LSD-0. Our results show that BP+LSD-0 achieves comparable decoding performance to the BP+OSD-CS-7 decoding results presented in Ref. [12] (where BP+OSD-CS-7 refers to the “combination sweep” strategy for OSD higher-order processing with order $w=7$, see Ref. [21]). This implies that, at least for the case of the BB code family, higher-order reprocessing has minimal impact when decoding in the sub-threshold regime. See Appendix˜D for further discussion of higher-order reprocessing.
Another interesting observation from our numerical decoding simulations of BB codes is that it is sufficient to run BP+LSD with a low number of BP iterations: the BP+LSD simulations shown in Fig. 6 of the main text were run with maximum $30$ BP iterations, compared to the maximum number of iterations of $10,000$ for the BP+OSD-CS-7 simulations in Ref. [12]. For BP-OSD, it is preferable to run BP until convergence to avoid the cost of computing the inverse of the detector check matrix. In OSD, matrix inversion is performed globally with worst-case cubic runtime complexity in the number of fault nodes in the Tanner graph $n$. Therefore, the OSD runtime does not scale with the error rate and is independent of the number of flipped detectors. In contrast, this is not the case for our proposed BP+LSD decoder. When run in parallel, the LSD algorithm has an expected runtime in $O(\kappa^{3})$, where $\kappa$ is the maximum cluster size. Further to this, the number of clusters upon initialization is identical to the number of invalid detectors. We therefore expect the decoding time to be proportional to the physical error rate. As such, in the sub-threshold regime, LSD post-processing is much less costly than BP+OSD. As a result, it is not detrimental to the decoder’s runtime if a large portion of the computational load is shouldered by the LSD post-processing.
@@
C.4 Runtime estimation
In this section, we conduct preliminary simulations to investigate the runtime of the proposed prototypical open-source implementation of the decoder available at [73]. We would like to highlight that we leave an in-depth investigation of a fully parallel and optimized implementation as future work. The results presented here should be seen as a mere representation of the runtime of the currently available software implementation acting as a benchmark for future developments, that is, in particular, it is not meant as a real-time decoder.
Fig. 8 shows timing statistics for BP+LSD and BP+OSD. The violin plots depict the distributions of the timings of BP+OSD and BP+LSD, where circles represent the means. This data is post-selected on cases where LSD/OSD is called, i.e., where BP does not converge, such that one measures more accurately the performance of the post-processing routines. To highlight the overhead introduced by the post-processor, triangles depict the maximum runtime of BP on shots for which it converged, presenting a lower bound on the achievable timing statistics. For the BP decoder, we have used a parallel schedule and min-sum update rules.
From the plot, it is evident that for the current software implementation of BP+LSD, the LSD step only marginally increases the overall runtime, as opposed to for BP+OSD, where the OSD post-processing time induces significant runtime overhead compared to the runtime of BP alone.
Figure 8: Timing statistics for circuit-level noise decoding of the $\llbracket 144,12,12\rrbracket$ bivariate bicycle code at a physical error rate of $p=0.1\%$. The timing results are post-selected on BP non-convergence. Triangle markers indicate the longest time BP took if it converged.
To estimate the runtime behaviour of the decoder in the “realistic scenario”, for instance, in a numerical experiment, we also investigate timing data that is not post-selected on the BP-non convergence. Fig. 9 depicts the total timing statistics of a BP+LSD/OSD call, regardless of BP convergence, i.e., for some samples, BP might converge and LSD/OSD is not called.
Figure 9: Runtime for the decoding of the $\llbracket 144,12,12\rrbracket$ BB code instance at physical noise rate $p=0.1\%$ (top left), $p=0.2\%$ (top right), $p=0.3\%$ (center left), $p=0.5\%$ (center right), and $p=0.7\%$ (bottom left), that is, above the pseudo-threshold of the code [12]. Circle markers show average decoding time per shot.
We observe that with an increased number of BP iterations, the decoding times decrease, perhaps due to more BP convergence and thus fewer post-processor calls. The distributions of BP+OSD exhibit a bimodal form, corresponding to the cases where BP converges or does not. For a larger number of BP iterations, this effect diminished, again perhaps due to the increased convergence rate of BP. This is not the case for BP+LSD, whose distributions do not exhibit this bimodality form because the overhead induced by LSD is (on average) comparable to BP (and lower than OSD), but its distributions feature long tails.
@@ Appendix D Higher-order LSD
The central idea of higher-order LSD is to add an additional reprocessing step to LSD0-0 to achieve better decoding accuracy. Once LSD0-0 has terminated, meaning that all clusters are valid and local solutions have been computed, we conduct, up to $\mu$ additional growth step on each valid cluster. On the grown clusters, we then apply standard OSD reprocessing: we do this separately on each cluster matrix and therefore we preserve LSD locality. We note that: